How many ways can you fill three different positions by choosing from 15 different people? I keep getting 15 3.
Answer by Alan3354(69443)

How many ways can you fill three different positions by choosing from 15 different people? I keep getting 15 3.
----------------
For the 1st choice, you have one of 15. The 2nd is one of 14, the 3rd one of 13.
So it's 15*14*13 = 2730
But, if you choose any of the people 1st, 2nd or 3rd, it's the same group (called without ordering), so you have to divide by 3*2*1 (= 6).
2730/6 = 455.